# Optimizing Data Transfer Performance and Energy Efficiency with Deep Reinforcement Learning

### v1

@@ 3.2.3 Adaptive DRL Framework and Interrelation of Objectives
$r_{t}=\begin{cases}U(T_{t},L_{t}),&\text{(Fairness \& Efficiency)}\\[5.69054pt] \frac{T_{t}}{E_{t}},&\text{(Throughput-Focused Energy)}\end{cases}$ (6)
@@
3.3 DRL State Space, Actions and Reward
3.3.1 DRL State Space
Designing an effective state space requires focusing on features that reliably indicate network conditions without directly including the performance metrics (e.g., throughput, energy) the agent aims to optimize. This separation ensures the agent learns the relationship between its actions and these metrics indirectly, yielding more robust, generalizable policies. Rather than using raw, highly variable measurements, we extract more stable indicators at each Monitoring Interval (MI), for instance:
  • $RTT$ gradient: Reflects how $RTT$ changes over time rather than its instantaneous value.
  • $RTT$ ratio [48]: Compares the current mean $RTT$ to the minimum observed mean $RTT$ since the start of the session, normalizing against the session’s best performance.
  • Packet loss rate ($plr$): Highlights network congestion more directly than throughput or energy values, which are the final optimization targets.
The agent’s controllable settings—concurrency ($cc$) and parallelism ($p$)—are also included in the state, enabling the RL algorithm to learn how past parameter choices affect the current network state.
To capture temporal dynamics, we store $n$ consecutive observations (one per MI). Let $x_{t}$ be the signal vector at time $t$:
$x_{t}=\{plr_{t},\,rtt\_gradient_{t},\,rtt\_ratio_{t},\,cc_{t},\,p_{t}\}.$ (7)
Then the state $s_{t}$ is:
$s_{t}=(x_{t-n},\ldots,x_{t-1},x_{t}),$ (8)
allowing the agent to detect trends in $plr$, $RTT$ metrics, and parameter updates over the last $n$ intervals.
By tracking these features over time, the agent learns how past adjustments to $(cc,p)$ influence future network conditions, discovering how to indirectly maximize throughput and minimize energy consumption. In essence, this state design balances simplicity and expressiveness, helping the RL approach adapt effectively to real-world, dynamically changing networks.
3.3.2 Action Space and Parameter Constraints
All the DRL algorithms we evaluate – on-policy (PPO [49], R_PPO [50]) and off-policy (DDPG [51], DQN [52], DRQN [53]) – use a discrete action space of five actions that jointly modify concurrency ($cc$) and parallelism ($p$). At each monitoring interval (MI) $t$, the agent picks an action $a_{t}\in\{0,1,2,3,4\}$, which updates $(cc_{t+1},p_{t+1})$ as follows:
$a_{t}=0\;\;$ $\rightarrow\;(cc_{t+1},p_{t+1})=(cc_{t},\,p_{t})$
$a_{t}=1\;\;$ $\rightarrow\;(cc_{t+1},p_{t+1})=(cc_{t}+1,\,p_{t}+1)$
$a_{t}=2\;\;$ $\rightarrow\;(cc_{t+1},p_{t+1})=(cc_{t}-1,\,p_{t}-1)$
$a_{t}=3\;\;$ $\rightarrow\;(cc_{t+1},p_{t+1})=(cc_{t}+2,\,p_{t}+2)$
$a_{t}=4\;\;$ $\rightarrow\;(cc_{t+1},p_{t+1})=(cc_{t}-2,\,p_{t}-2)$
In methods like PPO, R_PPO, and DDPG, the policy can internally produce separate real-valued outputs for $cc$ and $p$ (e.g., $(x_{1},x_{2})\in\mathbb{R}^{2}$), which are then floored or capped to map them into one of the five discrete actions. This approach supports finer control over concurrency and parallelism, yet still converges on the same discrete final choice.
We start with predefined $(cc_{0},p_{0})$ (e.g., $(4,4)$), derived from prior logs or experience, to avoid random extremes. Additionally, we enforce:
$cc_{\min}\leq cc_{t}\leq cc_{\max}\quad\text{and}\quad p_{\min}\leq p_{t}\leq p_{\max},$ (9)
clipping any actions that would exceed these limits. By bounding concurrency and parallelism in this way, we avoid overloading the network and end hosts while ensuring performance remains above a minimum baseline.
Although each DRL algorithm handles exploration differently (e.g., continuous outputs vs. discrete Q-tables), all converge on the same five discrete updates. This uniform action design simplifies comparisons across algorithms and ensures a clear mapping between each algorithm’s decisions and the concurrency ($cc$) and parallelism ($p$) configuration. Overall, this setup allows the agent to explore concurrency and parallelism in a controlled manner, preserving system stability and enabling effective parameter tuning.
3.3.3 Reward Formulation under Different Objectives
Our DRL agent’s reward function $r_{t}$ is designed to steer learning toward specific performance objectives. We primarily consider two reward formulations—one that promotes both fairness and efficiency (by reducing congestion and indirectly saving energy), and another that explicitly targets throughput per unit energy. This framework can be extended to additional objectives by tailoring the reward function.
Fairness and Efficiency Reward:
This reward is based on the utility function $U(T,L)$ (see Section 3.2), which rewards higher throughput $T$ while penalizing higher packet loss $L$. By discouraging congestion (i.e., high loss), this objective not only promotes fair bandwidth sharing but also indirectly reduces energy wasted on retransmissions and idle waiting:
$U(T,L)=\frac{T}{K^{(cc\times p)}}-T\cdot L\cdot B,$ (10)
For each measurement interval, the agent computes utility values $\{U_{t-n+1},\dots,U_{t}\}$ and forms an average:
$\overline{U}_{t}=\frac{1}{n}\sum_{i=t-n+1}^{t}U_{i}.$ (11)
The reward is then defined as:
$r_{t}=f\bigl{(}\overline{U}_{t},r_{t-1}\bigr{)},$ (12)
where $f(\cdot)$ measures the improvement relative to a rolling baseline.
Throughput-Focused Energy Efficiency Reward:
For energy-sensitive environments, the reward is defined based on throughput achieved per unit energy consumed. During each MI, let $T_{i}$ denote the throughput and $E_{i}$ the energy consumption. We compute:
$\overline{T}_{t}$ $=\frac{1}{n}\sum_{i=t-n+1}^{t}T_{i},\quad\text{and}\quad\overline{E}_{t}=\max_{i\in[t-n+1,t]}E_{i},$ (13)
which yields an energy efficiency metric:
$\overline{R}_{t}=\frac{\overline{T}_{t}\cdot SC}{\overline{E}_{t}},$ (14)
where $SC$ is a scaling constant. The corresponding reward is given by:
$r_{t}=f\bigl{(}\overline{R}_{t},r_{t-1}\bigr{)}.$ (15)
This formulation explicitly incentivizes higher throughput per unit energy, aligning with energy-constrained operation.
Difference-Based Reward Update ($f(\cdot)$):
A practical choice for $f(\cdot)$ leverages the difference between consecutive rewards to encourage incremental improvements. Specifically, let $r_{t}$ be the current metric (either $\overline{U}_{t}$ or $\overline{R}_{t}$) and $r_{t-1}$ be the previous one. We define:
  • $f(r_{t},r_{t-1})=x$, if $r_{t}-r_{t-1}>\epsilon$,
  • $f(r_{t},r_{t-1})=y$, if $r_{t}-r_{t-1}<-\epsilon$,
  • $f(r_{t},r_{t-1})=0$, otherwise,
where $x$ and $y$ are positive and negative rewards, respectively, and $\epsilon$ controls sensitivity. If $\epsilon$ is too small, the agent may oscillate due to frequent reward changes; if it is too large, learning may slow as only substantial improvements trigger non-zero rewards.
3.3.4 Unified DRL Optimization Problem
Regardless of the chosen DRL algorithm (PPO, R_PPO, DDPG, DQN, or DRQN) and the specific action space configuration, the agent aims to maximize the expected sum of discounted rewards:
[…]
where $\gamma\in(0,1]$ is the discount factor and $r_{t}$ is defined according to the selected performance objective. On-policy methods (PPO, R_PPO) estimate the policy gradient using samples drawn from the current policy, while off-policy methods (DDPG, DQN, DRQN) leverage replay buffers or target networks. The selected action space representation and reward function define the agent’s experience and, ultimately, its learned policy.
By appropriately choosing the action space structure and the reward function, we ensure that the agent can adapt to different optimization goals (fairness or energy efficiency), handle various RL algorithms, and dynamically learn to set optimal concurrency and parallelism parameters.
Figure 2: Overview of the offline-online training process, which involves a real environment where data is collected from the source (Src) through the network (Net) to the destination (Dst). The collected data undergoes exploration and logging, followed by clustering using k-means. The resulting data is used for training DRL agents in an emulated environment, with a feedback loop for validation and re-training of the policy.
3.4 Emulated Training Environment
To accelerate the DRL agent’s learning and minimize the time and energy costs associated with suboptimal data transfers, we built an emulated training environment from previously recorded state-action transitions and instantaneous transfer metrics. Traditional online training requires continuous interactions with a real system, causing many low-efficiency transfers before reaching a near-optimal policy. In contrast, our emulated setup refines the agent’s decision-making offline, reusing previously observed transitions to approximate real-world dynamics without repeatedly incurring inefficient runs. Figure 2 summarizes this offline-online process.
First, we deploy the DRL agent in a real environment under a high-exploration regime. During this phase, each second of the Monitoring Interval (MI) generates a transition log, recording instantaneous metrics such as throughput, loss rate ($plr$), round-trip time (RTT), energy consumption, and current $(cc,p)$. Each entry also includes a utility score derived from the chosen reward function. An illustrative entry might be:
  • 1707718539.468927 -- INFO:
    Throughput:8.32Gbps
    lossRate:0
    parallelism:7
    concurrency:7
    score:3.0
    rtt:34.6ms
    energy:80.0J
[…]
To structure these transitions, we convert each second’s measurements (e.g., throughput, RTT, $plr$, $(cc,p)$) into a feature vector:
$x=[plr,\;rtt\_gradient,\;rtt\_ratio,\;cc,\;p]$ (17)
Hence, each transition becomes $(x_{t},a_{t},x_{t+1},U_{t})$, where $U_{t}$ denotes the utility score. We then apply a clustering algorithm, such as $k$-means [54], grouping transitions that share similar features and outcomes. Each cluster’s centroid represents a recurring “network scenario,” enabling a simplified lookup when simulating new training episodes.
Once clustered, our emulator acts as a lookup-based environment:
@@ 3.4 Emulated Training Environment
    Initialization: Randomly pick an initial state $(cc_{0},p_{0})$ from the dataset for the start of a training episode.
@@
    Action Selection: Whenever the DRL agent chooses an action $a_{t}$, identify the cluster whose centroid most closely matches $(x_{t},a_{t})$.
  3. 3.
    Uniform Sampling: Randomly select a transition from the matching cluster to produce $x_{t+1}$ and the associated throughput, loss, and energy. This step simulates what might happen in a real transfer, without performing another physical run.
Each cluster contains multiple similar transitions, introducing beneficial variability into the next state and preventing overfitting to a single deterministic mapping.
[…]
Once the agent demonstrates strong performance in the emulator, we validate and, if necessary, re-train it in the real environment (see Figure 2). This iterative offline-online process keeps training costs low while preserving real-world accuracy; any newly gathered transitions can be clustered and added to the emulator, ensuring it remains up-to-date with evolving network states.
Figure 3: Diagram illustrates the primary DRL components for optimizing transfer parameters. The agent observes the state features (e.g., $plr,rtt\_gradient,rtt\_ratio,cc,p$), takes an action ($\{\Delta cc,\Delta p\}$), and receives a reward based on fairness or energy-efficiency objectives, all within an environment.
3.5 Evaluated DRL Algorithms
In this work, we consider a variety of DRL algorithms to handle the complex problem of optimizing concurrency ($cc$) and parallelism ($p$) under varying network conditions. Figure 3 illustrates how the agent observes recent states and selects actions to adjust $cc$ and $p$, receiving rewards (e.g., fairness or energy efficiency) from the environment. We evaluate both on-policy and off-policy algorithms to ensure that our approach is robust across different training paradigms. In particular, the five algorithms—DQN, DRQN, PPO, R_PPO, and DDPG—were chosen as representative methods for their unique strengths: DQN for its simplicity; DRQN and R_PPO for handling the partially observable network state via recurrent structures that capture patterns from state history; and PPO and DDPG as examples of on-policy and off-policy approaches, respectively, covering a broad spectrum of reinforcement learning strategies.
Deep Q-Networks (DQN) [52] combine Q-learning with deep neural networks, enabling the agent to approximate Q-values for high-dimensional state spaces. We construct a state vector by concatenating a fixed number of recent state observations, allowing the network to capture short-term temporal patterns. While effective for many tasks, DQN struggles with continuous actions and partially observable environments. Its simplicity makes it a useful baseline for our evaluation.
Proximal Policy Optimization (PPO)  [49] is a policy gradient method that adjusts the policy parameters to improve performance while constraining large updates. Similar to DQN, we represent the state as a concatenation of recent observations. PPO’s clipping mechanism ensures stable and efficient learning, making it suitable for continuous or discrete actions. As an on-policy method, it complements the off-policy approaches and provides insights into the benefits of policy gradient methods for our problem.
Deep Deterministic Policy Gradient (DDPG)  [51] is designed for continuous action spaces. It uses actor-critic architectures, where the actor outputs deterministic actions and the critic estimates Q-values. As with DQN and PPO, we feed a concatenated state vector to both actor and critic networks. DDPG represents the off-policy continuous control paradigm, making it an important candidate for optimizing the fine-grained adjustments required in network tuning.
Recurrent PPO (R_PPO) extends PPO by incorporating recurrent neural network layers (e.g., LSTMs) to handle partial observability and longer-term dependencies in the data [55, 50]. Instead of concatenating multiple states into a single vector, a window of past states is passed through the recurrent structure, enabling the agent to maintain an internal memory of historical network conditions and better adapt to fluctuating environments. This recurrent formulation is crucial for capturing the temporal patterns in network performance metrics.
Deep Recurrent Q-Networks (DRQN)  [53] augment DQN with recurrent units, allowing the agent to handle partially observable scenarios by storing information over time. Similar to R_PPO, the DRQN agent receives a window of states, feeding them sequentially into its recurrent layer. This approach provides temporal context and improves decision-making when immediate observations are insufficient.
By evaluating these algorithms, we determine which DRL method best captures the temporal dynamics, handles partial observability, and effectively optimizes throughput and energy consumption. The inclusion of both on-policy and off-policy methods, as well as models with and without recurrent architectures, ensures a comprehensive assessment of DRL techniques.
10:  Models: A chosen DRL model from {DQN, DRQN, PPO, R_PPO, DDPG};
@@ 3.5 Evaluated DRL Algorithms
42:  Define maximum number of training episodes $N$ and per-episode length $T$
3:  for episode $=1$ to $N$ do
@@
6:        Observe current state $s_{t}$ (e.g., $plr$, $rtt\_gradient$, $rtt\_ratio$, $cc$, $p$)
7:        Select action $a_{t}$ according to the DRL method:
@@
516:  return  policy $\pi_{\theta}$ (and $Q_{\psi}$ or $V_{\psi}$ if applicable).
@@
Table 1: Comparison of RL algorithms trained for 100000 steps with associated training metrics.
This high-level procedure (inspired by the reinforcement learning frameworks described in algorithm  1 outlines how various DRL models are trained to adjust concurrency ($cc$) and parallelism ($p$) in dynamic network environments.
  • Lines 1–2: Initialization. The agent and environment are set up according to the chosen RL model (DQN, DRQN, PPO, R_PPO, or DDPG). The model parameters (e.g., neural network weights) are randomly initialized, and an experience buffer (for off-policy methods) is prepared.
  • Lines 3–4: Outer Training Loop. The algorithm runs for a predefined number of episodes $N$, resetting the environment at the start of each episode.
  • Lines 5–10: Per-Step Operations. At each time step, the agent observes the current state (e.g., history of $plr$, $rtt\_gradient$, $rtt\_ratio$, $cc$, $p$), selects an action (increment or decrement $cc$ and $p$), applies it to the environment, and receives a reward reflecting the objective (fairness or energy efficiency).
  • Lines 11–12: Storage and Update. The agent stores the transition $(s_{t},a_{t},r_{t},s_{t+1})$ in either a replay buffer (off-policy) or a trajectory buffer (on-policy) and updates its policy or Q-value function using one of the RL optimization methods. This step leverages the well-known gradient-based updates [52, 49, 51] to minimize loss functions or maximize clipped surrogate objectives.
  • Lines 13–14: Termination Condition. Each episode ends when the agent reaches the final time step $T$ or when an environment-defined terminal state occurs (e.g., completion of a transfer or a set of transfer intervals).
  • Final Return. After $N$ episodes, the algorithm produces a trained policy $\pi_{\theta}$ (and possibly a learned value function $V_{\psi}$ or Q-function $Q_{\psi}$), which can then be used to select $cc$ and $p$ in real or emulated network environments.
3.6 Comparison of DRL Solutions
This section presents the comparison of five DRL algorithms (DQN, PPO, DDPG, R_PPO, and DRQN) for throughput focused energy objective (T/E) of data transfer optimization. We begin with a brief overview of the setup for offline training, followed by an analysis of the metrics in Table 1. Finally, we discuss the throughput and energy performance in both simulation and real-world transfers.
3.6.1 Offline Training Setup
To facilitate offline training, we first gathered transition data from a real-world environment (Chameleon Cloud, TACC to UC), with a wide range of $cc$ and $p$ settings. These transitions, logged at per-second intervals, captured throughput, packet loss rate ($plr$), round-trip time (RTT), and energy usage. By including both suboptimal and near-optimal parameter combinations, the dataset covered diverse network conditions (e.g., congestion levels, and background load). Each transition comprised a state (featuring $plr$, $rtt\_gradient$, $rtt\_ratio$, $cc$, $p$), an action (changes to $cc$ and $p$), and the resulting throughput and energy. We then employed clustering-based emulation (Section 3.4) to train each DRL model without repeatedly interacting with the real network, achieving more efficient policy convergence.
3.6.2 Analysis of Results in Table 1
Table 1 compares five DRL algorithms across several metrics:
Offline Training Time (min) $\downarrow$ and Steps in Converging $\downarrow$: These columns show the duration (wall-clock minutes) and interaction steps required for each method to reach a stable policy in offline mode. For example, DQN converges fastest (30 minutes, 300K steps), while DRQN is slowest (94 minutes), likely due to its recurrent architecture and additional computation needed to handle partial observability.

### v2

@@ III-B3 Unifying View
$r_{t}=\begin{cases}U_{\mathrm{FE}}(T_{t},L_{t},n_{t}),&\text{(Fairness \& Efficiency)},\\[2.84526pt] U_{\mathrm{TE}}(T_{t},E_{t}),&\text{(Throughput/Energy)}.\end{cases}$ (8)
@@
III-C Theoretical Validation: Solvability and Utility Geometry
III-C1 Existence of an Optimum
We consider a quasi-stationary network condition $\omega$ (e.g., cross-traffic intensity, queueing state, end-host contention) over a short monitoring interval (MI). For any choice of concurrency $cc$ and parallelism $p$ within predefined safe bounds, the transfer produces measurable outcomes: throughput $T(cc,p;\omega)$, loss $L(cc,p;\omega)$, and energy $E(cc,p;\omega)$. We denote the total number of TCP streams by
$n\triangleq cc\cdot p,\qquad(cc,p)\in\mathcal{A},$ (9)
where $\mathcal{A}$ is a finite feasible set (bounded by $cc_{\min}\!\leq cc\!\leq cc_{\max}$ and $p_{\min}\!\leq p\!\leq p_{\max}$).
For a fixed $\omega$, the instantaneous utility (either fairness & efficiency or throughput/energy) is therefore a function over a finite action set:
$(cc^{*}(\omega),p^{*}(\omega))\in\arg\max_{(cc,p)\in\mathcal{A}}U(cc,p;\omega).$ (10)
Hence, the static optimization is solvable. The practical difficulty is not existence, but that $U(\cdot;\omega)$ is unknown a priori and the underlying condition $\omega$ changes during the transfer, so the maximizer drifts over time.
III-C2 Why a Dominant Optimum is Expected
For loss-based TCP on wide-area paths, increasing the number of streams typically yields (i) diminishing throughput gains near the bottleneck capacity and (ii) sharply increasing congestion signals (loss and RTT inflation) once the bottleneck is persistently overfilled. Combined with stream-management overhead, this produces a “knee”: below the knee, adding streams helps; beyond it, congestion and overhead dominate.
Formally, for a quasi-stationary $\omega$, the following mild behaviors are commonly observed:
  • $T(cc,p;\omega)$ increases with diminishing returns as the bottleneck saturates,
  • $L(cc,p;\omega)$ stays low initially and rises rapidly once persistent queueing begins,
  • overhead increases with the stream budget $n=cc\cdot p$ (e.g., via the $K^{n}$ term in Eq. (4)).
These effects yield a utility landscape with a dominant peak around the knee, i.e., a unimodal (quasi-concave) structure over the feasible $(cc,p)$ grid for a fixed $\omega$.
Figure 2(a) abd  2(b) visualizes exactly this: the measured utility over $(cc,p)$ exhibits one clear maximizer (marked in the plot) and decreases smoothly away from that point, indicating that aggressively increasing streams eventually reduces utility due to congestion and overhead. Figure 3 further supports trackability: starting from diverse initial configurations, a local update rule converges to the same maximizer within a small number of steps. This behavior motivates our bounded, small-step action design ($\pm 1,\pm 2$), which is sufficient to reach (and later track) the moving optimum as $\omega_{t}$ changes.
Figure 3: Local-search trajectories. Local updates from diverse initializations converge to the same maximizer, suggesting small bounded steps are sufficient to reach (and track) the optimum.
III-D RL Formulation (POMDP)
Because the true network condition $\omega_{t}$ is not directly observable at end hosts, we model tuning as a Partially Observable Markov Decision Process (POMDP). At each MI $t$, the agent observes an end-host feature history $s_{t}$, selects an action $a_{t}$ that updates $(cc_{t},p_{t})$, and receives reward $r_{t}$ defined in Eq. (8). The RL objective is:
$\max_{\theta}\;\mathbb{E}_{\pi_{\theta}}\!\left[\sum_{t=t_{s}}^{t_{f}}\gamma^{t-t_{s}}r_{t}\right],\qquad\gamma\in(0,1],$ (11)
where $\pi_{\theta}(a_{t}\mid s_{t})$ is the learned policy parameterized by $\theta$. Intuitively, the agent learns to increase $n_{t}$ when the path is underutilized and to decrease $n_{t}$ when loss/RTT trends indicate rising congestion, thereby tracking the moving optimum implied by the unimodal utility geometry.
III-E State, Action and Reward Design
III-E1 RL State Space
The agent must infer congestion from end-host signals, so the state uses stable indicators rather than the optimization targets (throughput/energy). At each monitoring interval (MI) $t$, we form:
$x_{t}=\{plr_{t},\ rtt\_gradient_{t},\ rtt\_ratio_{t},\ cc_{t},\ p_{t}\},$ (12)
where $plr_{t}$ is packet loss rate, $rtt\_gradient_{t}$ captures RTT trend, and $rtt\_ratio_{t}$ normalizes current RTT by the best (minimum) RTT observed so far [31]. We include $(cc_{t},p_{t})$ so the policy can associate parameter choices with subsequent congestion signals. To mitigate partial observability and measurement noise, the agent observes a short history window:
$s_{t}=(x_{t-h+1},\ldots,x_{t}),$ (13)
which allows the policy to detect trends (e.g., rising RTT and loss) and react before severe congestion.
III-E2 Action Space and Parameter Constraints
All evaluated RL methods (PPO [53], R_PPO [32], DDPG [38], DQN [45], DRQN [26]) use the same discrete action set of five joint updates to $(cc,p)$. At MI $t$, the agent selects $a_{t}\in\{0,1,2,3,4\}$:
$a_{t}=0$ $:(cc_{t+1},p_{t+1})=(cc_{t},p_{t})$
$a_{t}=1$ $:(cc_{t+1},p_{t+1})=(cc_{t}+1,p_{t}+1)$
$a_{t}=2$ $:(cc_{t+1},p_{t+1})=(cc_{t}-1,p_{t}-1)$
$a_{t}=3$ $:(cc_{t+1},p_{t+1})=(cc_{t}+2,p_{t}+2)$
$a_{t}=4$ $:(cc_{t+1},p_{t+1})=(cc_{t}-2,p_{t}-2).$
We clip updates to enforce operational bounds:
$cc_{\min}\leq cc_{t}\leq cc_{\max},\qquad p_{\min}\leq p_{t}\leq p_{\max},$ (14)
and initialize from a moderate starting point (e.g., $(cc_{0},p_{0})=(4,4)$) to avoid extreme exploration early in an episode. Using a shared, bounded action space ensures controlled exploration (stability on real networks) and enables fair comparison across RL algorithms.
III-E3 Reward Computation and Smoothing
We compute $r_{t}$ using Eq. (8). To reduce measurement noise, we optionally use a short moving average:
$\tilde{r}_{t}=\frac{1}{h}\sum_{i=t-h+1}^{t}r_{i},$ (15)
and train on $\tilde{r}_{t}$.
Figure 4: Offline–online workflow: exploratory data collection in the real environment, clustering of transitions, offline training in the emulator, and periodic validation / re-training with new logs.
III-F Offline–Online Training via an Emulated Environment
Training an RL controller purely online is expensive because early exploration produces many suboptimal transfers. To reduce time and energy overhead, we train primarily offline using an emulator built from previously recorded transitions, and then validate (and optionally fine-tune) the policy online. Figure 4 summarizes this workflow.
Step 1: High-exploration data collection (real path)
We first run the agent on the real network under a high-exploration policy and log per-MI (or per-second) measurements, including throughput, loss rate (plrplr), RTT, energy, and the current $(cc,p)$. Each log entry is labeled with the reward/utility value produced by the chosen objective. Conceptually, this yields a dataset of transitions
$\mathcal{D}=\{(s_{t},a_{t},r_{t},s_{t+1})\},$ (16)
where $s_{t}$ is the observed state (Section III-E1), $a_{t}$ is the discrete action (Section III-E2), and $r_{t}$ is the reward.
Step 2: Clustering transitions into “network scenarios”
We map each MI to a compact feature vector (e.g., $x_{t}=[plr_{t},\,rtt\_gradient_{t},\,rtt\_ratio_{t},\,cc_{t},\,p_{t}]$) and cluster transitions using $k$-means with $k=40$ clusters [39]. Each cluster groups similar state–action contexts and their observed outcomes, and its centroid represents a recurring ”network scenario,” enabling a simplified lookup when simulating new training episodes.
Figure 5: RL loop for transfer tuning: the agent observes state features (history of plrplr, RTT trends, and current $(cc,p)$), selects an action ($\Delta cc,\Delta p$), and receives reward consistent with the selected objective.
Step 3: Lookup-based emulation for offline training
The emulator behaves like an environment that returns plausible next states and rewards without executing a real transfer:
@@ III-F Offline–Online Training via an Emulated Environment
    Initialize: sample an initial state $s_{1}$ from $\mathcal{D}$.
@@
    Step: given $(s_{t},a_{t})$, select the nearest cluster (or clusters) and uniformly sample one transition from that cluster to produce $(s_{t+1},r_{t})$.
Sampling within a cluster introduces variability (multiple plausible outcomes) and reduces overfitting compared to a deterministic lookup. Figure 6 shows that trends learned offline in the emulator transfer to real deployments: algorithms that achieve higher throughput (or lower energy) in emulation exhibit consistent behavior during real transfers, supporting the emulator’s fidelity for offline training.
Step 4: Online validation and refresh.
After the policy converges offline, we deploy it on the real environment for validation. If performance degrades due to a new network regime, we can collect additional transitions, re-cluster, and refresh the emulator or retune the policy in real world environment.
III-G Evaluated RL Algorithms and Training Procedure
We evaluate representative on-policy and off-policy RL methods for tuning $(cc,p)$: DQN [45], DRQN [26], PPO [53], R_PPO [55] , and DDPG [38]. We include recurrent variants (DRQN, R_PPO) because end hosts only observe indirect congestion signals; recurrence provides memory that improves decisions under partial observability. Figure 5 shows the closed loop: the agent observes the recent state, selects an action to update $(cc,p)$, and receives a reward defined by the selected objective.
10:Models: A chosen RL model from {DQN, DRQN, PPO, R_PPO, DDPG};
@@ III-G Evaluated RL Algorithms and Training Procedure
42: Define maximum number of training episodes NN and per-episode length $T$
3:for episode =1=1 to NN do
@@
6:   Observe current state $s_{t}$ (e.g., plrplr, rtt_gradientrtt\_gradient, rtt_ratiortt\_ratio, cccc, $p$)
7:   Select action $a_{t}$ according to the RL method:
@@
16:return policy $\pi_{\theta}$ (and $Q_{\psi}$ or $V_{\psi}$ if applicable).
@@
Algorithm overview. Algorithm 1 follows a standard RL loop: initialize the model and (if applicable) a replay buffer; repeatedly interact with the environment by observing $s_{t}$, choosing $a_{t}$, receiving $(r_{t},s_{t+1})$, storing experience, and updating the model. On-policy methods (PPO, R_PPO) update from recent rollouts of the current policy, while off-policy methods (DQN, DRQN, DDPG) reuse stored transitions via a replay buffer for sample-efficient learning. The output is a trained policy that selects $(cc,p)$ online from end-host observations.
@@
TABLE I: Comparison of RL algorithms trained for 100000 steps with associated training metrics.
III-H Comparison of RL Solutions
We compare five RL algorithms (DQN, PPO, DDPG, R_PPO, DRQN) under the throughput-focused energy objective (T/E). We first describe the offline training data and emulator, then summarize training/inference costs (Table I), and finally compare transfer throughput and energy in simulation and on a real wide-area environment (Figure 6).
III-H1 Offline Training Setup